OpenStudy (anonymous):

Find the common difference of the arithmetic sequence. 4, 10, 16, 22, . . .

OpenStudy (kirbykirby):

10-4 = ? 16-10 = ? 22 - 16 = ?

OpenStudy (anonymous):

6

OpenStudy (kirbykirby):

yes